Find the liquid pressure at the bottom of a 50m de
Answer #1

21)


depth, h=50m


density of ethyl alcohol, rho=789 kg/m^3


pressure =rho*g*h


=789*9.8*50


=386610 Pascal

22)


temeperature is constant


initial pressure, P1=500 Pascal


volume, v1=10 m^3 , v2=2 m^3


use,


P1*v1=P2*v2


500*10=p2*2


===> p2=2500 Pascal


final pressure, p2=2500 Pascal


23)

mass of the sphere, m=25kg


area, A=10 m^2

pressure, P=F/A


P=m*g/A


P=(25*9.8)/(10)


P=24.5 Pascal